determine whether each integral is convergent or divergent

1- 1/(x-2)^3/2 dx, limits ( infinite to 3)

2- (1/3-4x)dx, limits (0 to -infinite)

3- e^(-5p) dx, limits ( infinite to 2)

4-   (x^2/(sqrt(1+x^3)))dx , limits ( infinite to 0)

5- lnx/x dx , limits(infinite to 1)

6- 1/(x^2 +x)dx , limits (infinite to 1)

7- 3/x^5 dx ,limits (1 to 0)

8- dx/(x+2)^1/4 , limits (14 to -2)

9- 1/(x-1)^1/3 , limits (9 to 0)

10- e^x /((e^x) -1), limits (1 to -1)

11- z^2 lnz dz ,limits( 2 to 0)

12- s={(x,y) | x>= 1,0 <= y <=e^-x}
Answer #1

Consider the integral (x-2)2 (x 2)2 -lim ar χ-2)2 --lin l (x-2 )--dx lim | (C-3)-3.1 = _ lim 2 limi- χ-2)2 = 2 | lim 1-1 -(t-

2 Finite value the integral f 1d is con dx is convergent 3 1 ”(x-2)? 2) Consider the integralch 3-4.x dr 3 4.x 3- 4x In(3-4x)

3) Consider the integral le--pdr lim e2) m(t- 2 Thus, the integral has infinite value. Therefore, the integral lePdx is diver

4) Consider the integral dr ax lim Let 1+x3=11 If x = 0 then u 1 ax=-lim ctr du = _ lim lim u 2diu

=--11m 2 lim| t2-1 Thus, the integral has infinite value. Therefore, the integral fe dx is divergent.